# One Brookings Drive St. Louis, MO 63130 USA

### v1

@@ 1 Introduction
for $p$ with no zeros in $\mathbb{D}^{d}$ and no factors in common with $\tilde{p}$ (the constant $a$ can be absorbed into the polynomials). We can also simply write the rational inner function as $\frac{\tilde{p}}{p}$ by reflecting $p$ at a degree higher than its multidegree. In this way we have a simple uniform way of writing rational inner functions. We shall give essentially Pfister’s proof Pfister in the next section. With Theorem 1.1, it is easy to construct examples of rational inner functions (that is, if one can easily construct polynomials with no zeros on $\mathbb{D}^{d}$.)
@@
$\lim_{\theta\to 0}f_{0}(e^{i\theta},e^{-i\theta+ia\theta^{2}})=\lim_{\theta\to 0}\frac{ia\theta+\theta^{2}+O(\theta^{3})}{-ia\theta+\theta^{2}+O(\theta^{3})}=\frac{1+ia}{1-ia}$
@@ 2 Basic facts
Proof (Proof of Theorem 1.1)
@@
$\phi(z)=P(z)+z^{N}\frac{1-\tilde{P}P}{1+z^{N}\tilde{P}}$
@@ 3 Interpolation
This theorem can be addressed entirely using rational inner functions (in this case, finite Blaschke products). Indeed, it is possible to show (1) implies that we have a finite Blaschke product that interpolates, and conversely, by the approximation theorem, Theorem 2, we only need to verify necessity for rational inner functions.
@@
Theorem 3.1 (Agler’s Pick interpolation theorem)
@@
For now, let us remark that currently there is no known interpolation theorem for $\mathbb{D}^{d}$ when $d>2$. An ‘interpolation theorem’ would mean that we reduce interpolation for bounded analytic functions to searching over a finite dimensional set. Theorem 3.1 does this even if the conditions given could be difficult to check. The following is not even known:
@@ 4 Sums of squares
Theorem 4.1
Let $\phi(z,w)=\frac{\tilde{p}(z,w)}{p(z,w)}$ be a rational inner function, where $p\in\mathbb{C}[z,w]$ has no zeros in $\mathbb{D}^{2}$, no factors in common with $\tilde{p}(z,w)$, and $\tilde{p}$ has bidegree $(n,m)$. Then, there exist polynomials $E_{1},\dots,E_{n},F_{1},\dots,F_{m}\in\mathbb{C}[z,w]$ such that
@@
Returning to Theorem 4.1, it is worth pointing out that such sums of squares formulas are in general not unique. A trivial way they are not unique is that we could replace
@@
Theorem 4.2
Let $p\in\mathbb{C}[z,w]$ have no zeros in $\mathbb{D}^{2}$, no factors in common with $\tilde{p}$, and $\deg\tilde{p}=(n,m)$. Consider the sums of squares terms $\mathcal{E}(z,w)=\sum_{j=1}^{n}|E_{j}(z,w)|^{2}$, $\mathcal{F}(z,w)=\sum_{j=1}^{m}|F_{j}(z,w)|^{2}$ appearing in Theorem 4.1. Then, the following are equivalent
@@
It is probably not feasible to characterize all sums of squares decompositions, but it turns out it is possible to characterize the decompositions where the number of squares, $n$ for $\mathcal{E}$ and $m$ for $\mathcal{F}$, are minimal as in Theorem 4.1. The characterization is too technical to state here (see Theorem 9.3 of gKintreg ) but the most interesting aspect is that the minimal decompositions are in one-to-one correspondence with the invariant subspaces of a special pair of commuting contractive matrices.
Assume $p$ is given as in Theorem 4.1. Let
@@ 5 Boundary behavior
The key local result needed for proving Theorem 5.1 is the following:
Theorem 5.2
@@
Theorem 5.4 (Theorem 2.16 in BKPS )
@@
is bounded above and below near 00 in $\mathbb{H}^{2}$. Thus, for many problems we can replace $p$ with the simpler polynomial $[p]$. At the same time, Theorem 5.4 makes it possible to construct polynomials with no zeros in $\mathbb{H}^{2}$ or $\mathbb{D}^{2}$ with non-trivial Puiseux structure at a boundary point—something that was previously not known.
Theorem 5.4 also led to the following conjecture and now theorem characterizing rational functions bounded near a boundary point.
Theorem 5.5 (Theorem 5.2 of BKPS , Theorem 3 of kollar )
Assume $p$ as in Theorem 5.4. Given $q\in\mathbb{C}[x,y]$, the rational function $q/p$ is bounded in a neighborhood of $(0,0)$ intersected with $\mathbb{H}^{2}$ if and only if $q$ belongs to the product ideal
@@
Important prior work related to Theorems 5.4 and 5.5 was done in the papers BPS1 , BPS2 where the local geometry of the zero set of $p$ was connected to the local geometry of the level sets of rational inner functions now written in the upper half plane setting as $\phi(x,y)=\bar{p}(x,y)/p(x,y)$
@@
The data $2L_{1},\dots,2L_{k}$ from Theorem 5.4 can be used to measure the rate that the zero set of $p$ approaches $\mathbb{R}^{2}$ near $(0,0)$. Specifically, $K_{0}=\max\{2L_{1},\dots,2L_{k}\}$ is called the local contact order of $p$ at $(0,0)$ because
$\inf\{|\Im y|:p(x,y)=0\}\approx|x|^{K_{0}}.$
@@
Crucial to the approach of BPS1 , BPS2 was showing that the branches of the level sets (8) occur in bundles called horn regions. Figure 1 shows various level sets of a rational inner function. Note how they occur in two horn-shaped regions or bundles. These bundles are directly connected to the polynomials $q_{1}(x),\dots,q_{k}(x)$ in Theorem 5.4. Necessarily, for a rational inner function $\phi$, we have $\frac{\partial\phi}{\partial z}$ belonging to $L^{1}(\mathbb{T}^{2})$. This fact even extends to several variables. In two variables, the paper Pascoe-cara shows how to construct polynomials with given global contact order so that conditions in Theorem 5.6 can actually occur.
@@
Finally, in the preprint gKint we gave a complete characterization of when a rational function $Q/p$ (where $p$ has no zero in $\mathbb{H}^{2}$ and no factors in common with $\bar{p}$) is locally in $L^{\mathfrak{p}}$ in $\mathbb{R}^{2}$ near the origin. The main result involves checking the behavior of $Q$ along curves related to the polynomials $q_{1}(x),\dots,q_{k}(x)$ from Theorem 5.4. A few of the easiest to state offshoots are below. In the next theorem, we consider the space of $Q/p$ that are locally square integrable:
@@
Theorem 5.7
Assume $p$ satisfies the setup and conclusion of Theorem 5.4. Then,
@@
Another easy to state result related to Theorem 5.6 is the following.
Theorem 5.8 (Theorem 1.9 of gKint )
@@
It turns out that rational inner functions actually exhibit the ‘worst’ derivative integrability behavior among bounded rational functions.
@@
Given the data $2L_{1},\dots,2L_{k},M_{1},\dots,M_{k},q_{1}(x),\dots,q_{k}(x)$ from Theorem 5.4 can one always construct a polynomial $p\in\mathbb{C}[x,y]$ with no zeros in $\mathbb{D}^{2}$ associated to this data as in Theorem 5.4?
6 Higher dimensions
In this final section we look at what is known about rational inner functions in three or more variables. As indicated in Section 3, if one could understand interpolation for rational inner functions, it ought to be possible to understand interpolation more broadly but if the answer to Problem 1 is negative, then a different approach may be necessary.
It turns out that the natural generalizations of Agler’s Pick interpolation (Theorem 3.1) and the sums of squares theorem (Theorem 4.1) do not hold for three or more variables, and this is well-known to be connected to operator theoretic obstructions. An analytic function $f:\mathbb{D}^{d}\to\mathrm{d}$ satisfies a formula
@@ 6 Higher dimensions
The exact derivative integrability cut-offs are known for rational inner functions in two variables by Theorem 5.6. What are the possible integrability cut-offs in three or more variables? Specifically, given a rational inner function $\phi$ on $\mathbb{D}^{d}$ for $d>2$, what are all of the possible values of
@@
7 Acknowledgements
I would like to thank Alan Sola for offering comments on a draft of this paper (on extremely short notice!). Also, thanks to all the collaborators and friends who have contributed to the interesting (to me) topic of rational inner functions.

### v3

@@ 0.1 Introduction
for $p$ with no zeros in $\mathbb{D}^{d}$ and no factors in common with $\tilde{p}$ (the constant $a$ can be absorbed into the polynomials). We can also simply write the rational inner function as $\frac{\tilde{p}}{p}$ by reflecting $p$ at a degree higher than its multidegree. In this way we have a simple uniform way of writing rational inner functions. We shall give essentially Pfister’s proof Pfister in the next section. With Theorem 0.1.1, it is easy to construct examples of rational inner functions (that is, if one can easily construct polynomials with no zeros on $\mathbb{D}^{d}$.)
@@
$\lim_{\theta\to 0}f_{0}(e^{i\theta},e^{-i\theta+ia\theta^{2}})=\lim_{\theta\to 0}\frac{ia\theta^{2}+\theta^{2}+O(\theta^{3})}{-ia\theta^{2}+\theta^{2}+O(\theta^{3})}=\frac{1+ia}{1-ia}$
@@ 0.2 Basic facts
Proof (Proof of Theorem 0.1.1)
@@
$\phi(z)=P(z)+z^{N}\frac{z^{n}-\tilde{P}P}{1+z^{N}\tilde{P}}$
@@ 0.3 Interpolation
This theorem can be addressed entirely using rational inner functions (in this case, finite Blaschke products). Indeed, it is possible to show (1) implies that we have a finite Blaschke product that interpolates, and conversely, by the approximation result, Proposition 2, we only need to verify necessity for rational inner functions.
@@
Theorem 0.3.1 (Agler’s Pick interpolation theorem)
@@
For now, let us remark that currently there is no known interpolation theorem for $\mathbb{D}^{d}$ when $d>2$. An ‘interpolation theorem’ would mean that we reduce interpolation for bounded analytic functions to searching over a finite dimensional set. Theorem 0.3.1 does this even if the conditions given could be difficult to check. The following is not even known:
@@ 0.4 Sums of squares
Theorem 0.4.1
Let $\phi(z,w)=\frac{\tilde{p}(z,w)}{p(z,w)}$ be a rational inner function, where $p\in\mathbb{C}[z,w]$ has no zeros in $\mathbb{D}^{2}$, no factors in common with $\tilde{p}(z,w)$, and $\tilde{p}$ has bidegree $(n,m)$. Then, there exist polynomials $E_{1},\dots,E_{n}$ of bidegree at most $(n-1,m)$ and $F_{1},\dots,F_{m}$ of bidegree at most $(n,m-1)$ such that
@@
Returning to Theorem 0.4.1, it is worth pointing out that such sums of squares formulas are in general not unique. A trivial way they are not unique is that we could replace
@@
Theorem 0.4.2
Let $p\in\mathbb{C}[z,w]$ have no zeros in $\mathbb{D}^{2}$, no factors in common with $\tilde{p}$, and $\deg\tilde{p}=(n,m)$. Consider the sums of squares terms $\mathcal{E}(z,w)=\sum_{j=1}^{n}|E_{j}(z,w)|^{2}$, $\mathcal{F}(z,w)=\sum_{j=1}^{m}|F_{j}(z,w)|^{2}$ appearing in Theorem 0.4.1. Then, the following are equivalent
@@
It is probably not feasible to characterize all sums of squares decompositions, but it turns out it is possible to characterize the decompositions where the number of squares, $n$ for $\mathcal{E}$ and $m$ for $\mathcal{F}$, are minimal as in Theorem 0.4.1. The characterization is too technical to state here (see Theorem 9.3 of gKintreg ) but the most interesting aspect is that the minimal decompositions are in one-to-one correspondence with the invariant subspaces of a special pair of commuting contractive matrices.
Assume $p$ is given as in Theorem 0.4.1. Let
@@ 0.5 Boundary behavior
The key local result needed for proving Theorem 0.5.1 is the following:
Theorem 0.5.2
@@
Theorem 0.5.4 (Theorem 2.16 in BKPS )
@@
is bounded above and below near 00 in $\mathbb{H}^{2}$. Thus, for many problems we can replace $p$ with the simpler polynomial $[p]$. At the same time, Theorem 0.5.4 makes it possible to construct polynomials with no zeros in $\mathbb{H}^{2}$ or $\mathbb{D}^{2}$ with non-trivial Puiseux structure at a boundary point—something that was previously not known. The construction amounts to writing out a proposed local decomposition such as
$P(x,y)=(y+x+ix^{2}+x^{5/2})(y+x+ix^{2}-x^{5/2})=(y+x+ix^{2})-x^{5}$
which one can show has no zeros in the $\mathbb{H}^{2}\cap\mathbb{D}_{\epsilon}^{2}$. Then, we can find a pair of conformal maps $\phi_{1}(x),\phi_{2}(y)$ that map $\mathbb{H}$ to a disk in $\mathbb{H}$ that is tangent to $\mathbb{R}$ at 00 so that $Q=P(\phi_{1}(x),\phi_{2}(y))$ has no zeros in $\mathbb{H}^{2}$ yet we do not change the fact that a local factorization of $Q$ requires Puiseux series. We can clear the denominators of $Q$ to get a polynomial non-vanishing on $\mathbb{H}^{2}$ with non-trivial Puiseux structure at $(0,0)$. This process blurs the original data associated with $P$, so we pose it as a problem later (Problem 2) to show that we can construct global stable polynomials with prescribed local data.
Theorem 0.5.4 also led to the following conjecture and now theorem characterizing rational functions bounded near a boundary point.
Theorem 0.5.5 (Theorem 5.2 of BKPS , Theorem 3 of kollar )
Assume $p$ as in Theorem 0.5.4. Given $q\in\mathbb{C}[x,y]$, the rational function $q/p$ is bounded in a neighborhood of $(0,0)$ intersected with $\mathbb{H}^{2}$ if and only if $q$ belongs to the product ideal
@@
Important prior work related to Theorems 0.5.4 and 0.5.5 was done in the papers BPS1 , BPS2 where the local geometry of the zero set of $p$ was connected to the local geometry of the level sets of rational inner functions now written in the upper half plane setting as $\phi(x,y)=\bar{p}(x,y)/p(x,y)$
@@
The data $2L_{1},\dots,2L_{k}$ from Theorem 0.5.4 can be used to measure the rate that the zero set of $p$ approaches $\mathbb{R}^{2}$ near $(0,0)$. Specifically, $K_{0}=\max\{2L_{1},\dots,2L_{k}\}$ is called the local contact order of $p$ at $(0,0)$ because for real $x$ near 00 and $\epsilon>0$ sufficiently small
$\inf\{|\Im y|:p(x,y)=0,|y|<\epsilon\}\approx|x|^{K_{0}}.$
@@
Crucial to the approach of BPS1 , BPS2 was showing that the branches of the level sets (9) occur in bundles called horn regions. Figure 1 shows various level sets of a rational inner function. Note how they occur in two horn-shaped regions or bundles. These bundles are directly connected to the polynomials $q_{1}(x),\dots,q_{k}(x)$ in Theorem 0.5.4. Necessarily, for a rational inner function $\phi$, we have $\frac{\partial\phi}{\partial z}$ belonging to $L^{1}(\mathbb{T}^{2})$. This fact even extends to several variables. In two variables, the paper Pascoe-cara shows how to construct polynomials with given global contact order so that conditions in Theorem 0.5.6 can actually occur.
@@
Finally, in the preprint gKint we gave a complete characterization of when a rational function $Q/p$ (where $p$ has no zero in $\mathbb{H}^{2}$ and no factors in common with $\bar{p}$) is locally in $L^{\mathfrak{p}}$ in $\mathbb{R}^{2}$ near the origin. The main result involves checking the behavior of $Q$ along curves related to the polynomials $q_{1}(x),\dots,q_{k}(x)$ from Theorem 0.5.4. A few of the easiest to state offshoots are below. In the next theorem, we consider the space of $Q/p$ that are locally square integrable:
@@
Theorem 0.5.7
Assume $p$ satisfies the setup and conclusion of Theorem 0.5.4. Then,
@@
Another easy to state result related to Theorem 0.5.6 is the following.
Theorem 0.5.8 (Theorem 1.9 of gKint )
@@
Given the data $2L_{1},\dots,2L_{k},M_{1},\dots,M_{k},q_{1}(x),\dots,q_{k}(x)$ from Theorem 0.5.4 can one always construct a polynomial $p\in\mathbb{C}[x,y]$ with no zeros in $\mathbb{D}^{2}$ associated to this data as in Theorem 0.5.4?
0.6 Higher dimensions
In this final section we look at what is known about rational inner functions in three or more variables. As indicated in Section 0.3, if one could understand interpolation for rational inner functions, it ought to be possible to understand interpolation more broadly but if the answer to Problem 1 is negative, then a different approach may be necessary.
It turns out that the natural generalizations of Agler’s Pick interpolation (Theorem 0.3.1) and the sums of squares theorem (Theorem 0.4.1) do not hold for three or more variables, and this is well-known to be connected to operator theoretic obstructions. An analytic function $f:\mathbb{D}^{d}\to\mathrm{d}$ satisfies a formula
@@ 0.6 Higher dimensions
The exact derivative integrability cut-offs are known for rational inner functions in two variables by Theorem 0.5.6. What are the possible integrability cut-offs in three or more variables? Specifically, given a rational inner function $\phi$ on $\mathbb{D}^{d}$ for $d>2$, what are all of the possible values of
@@
0.7 Acknowledgements
I would like to thank Alan Sola for offering comments on a draft of this paper (on extremely short notice!). Also, thanks to all the collaborators and friends who have contributed to the interesting (to me) topic of rational inner functions. Finally, thank you to the anonymous referee for numerous helpful suggestions. An additional thanks goes to Rongwei Yang for pointing out a typo in the proof of Prop 2.